A hydrogen atom moving at a speed v absorbs a photon of wavelength 122 nm and stops. Find the value of v. Mass of a hydrogen atom = 1.67 x 10–27 kg.

Grade:12

1 Answers

Arun
6 years ago

energy of a photon is = hc/λ

 

visible light lies in 390-750 nm 

 

maximum energy for 390 nm

 

h=6.626 x  10-34 

 

c= 3 x 108

 

λ=390 x 10-9

 

 

 

E=5.09 x 10-19 Joules

 

 

Use the formula     lambda = h/mv

 

lambda = 127nm

 

h = 6.626*10-34 Js

 

mass (m)= 1.67*10-27 kg

 

 

 

Now evaluate do it as a practice ...

 

You get v =3.1 m/s

hope it clears